# Complexity of deep computations via topology of function spaces

### v3

@@ 2.3. Newton’s method as a CCS
Let $p(z)$ be a non-constant polynomial with complex coefficients. We say that $(L_{p},\mathcal{P},\Gamma_{p})$ is Newton’s method for $p(z)$ if:
  • $L_{p}$ is the set of all $z\in\mathbb{C}$ such that there exists an open neighborhood $U$ of $z$ such that every sequence in $\{N_{p}^{n}:n\in\mathbb{N}\}$ has a subsequence that converges uniformly on compact subsets of $U$, where
    $N_{p}(z)=z-\frac{p(z)}{p^{\prime}(z)}$
    and $N_{p}^{n}:=N_{p}\circ N_{p}\circ\cdots\circ N_{p}$ is the nth-iteration of $N_{p}$.
  • $\mathcal{P}:=\{P_{1},P_{2},P_{3}\}$ where
    $P_{1}(z)=\frac{2{\rm Re}(z)}{|z|^{2}+1},$
    $P_{2}(z)=\frac{2{\rm Im}(z)}{|z|^{2}+1},$
    $P_{3}(z)=\frac{|z|^{2}-1}{|z|^{2}+1}.$
  • $\Gamma_{p}:=\{N_{p}^{n}:n\in\mathbb{N}\}$.
Remarks 2.6.
  1. (1)
    The set $L_{p}$ is known as the Fatou set of $N_{p}$ (see section 2 in [BLA84]). Its complement $\mathbb{C}\backslash L_{p}$, is known as the Julia set of $N_{p}$.
  2. (2)
    The map $z\mapsto(P_{1}(z),P_{2}(z),P_{3}(z))$ is the stereographic projection into the Riemann Sphere $S^{2}$.
  3. (3)
    The set $L_{p}$ is open and dense in $\mathbb{C}$ ([BLA84], Corollary 4.6). Hence, its closure $\mathcal{L}$ in $\mathbb{R}^{\mathcal{P}}$ is the Riemann sphere $S^{2}$ (i.e., the extended complex plane).
  4. (4)
    The set $L_{p}$ is completely invariant under iterations of $N_{p}$, i.e., $N_{p}(L_{p})=L_{p}$ and $N_{p}^{-1}(L_{p})=L_{p}$. This implies that all iterations $N_{p}^{n}:L_{p}\rightarrow L_{p}$ are transition maps.
  5. (5)
    $\Gamma_{p}$ is the semigroup generated by $\{N_{p}\}$. Thus, $(L_{p},\mathcal{P},\Gamma_{p})$ is a CCS.
Newton’s method is an iterative method that is used to approximate a root of $p(z)$. The map $N_{p}(z)$ defined above is known as Newton’s map. The method consists of taking an initial guess $z_{0}\in\mathbb{C}$ and iterating the rational map $N_{p}$ to obtain a sequence given by
$z_{n+1}=N_{p}(z_{n})=z_{n}-\frac{p(z_{n})}{p^{\prime}(z_{n})}$
For each root $r\in\mathbb{C}$ of $p(z)$, there exists an $\varepsilon>0$ such that for any initial guess $z_{0}$ in the $\varepsilon$-ball centered at $r$, Newton’s iteration converges to $r$ (provided $p^{\prime}(r)\neq 0$) and the convergence is quadratic in that case, meaning the error at each step is roughly squared, causing the number of correct digits to double, leading to fast convergence.
Given a root $r$ of $p(z)$, the set $B_{r}=\{z\in\mathbb{C}:\lim_{n\rightarrow\infty}N_{p}^{n}(z)=r\}$ is an open set called the basin of $r$. However, Newton’s method can fail to converge to any root for some choices of $z_{0}$. For example, consider the polynomial $p(z)=z^{3}-2z+2$. The Newton map is given by
$N_{p}(z)=z-\frac{z^{3}-2z+2}{3z^{2}-2}=\frac{2z^{3}-2}{3z^{2}-2}$
Notice that taking $z_{0}=0$ as an initial guess will yield the sequence $0,1,0,1,0,1,\dots$ that oscillates between 0 and $1$ but none of them are roots of $p(z)$. Another more chaotic way Newton’s method can fail to converge is when the sequence of iterations has no convergent subsequence. The set of such points, i.e. the Julia set associated to $N_{p}$, is typically a fractal. This can be visualized by adding a dash of color: let us give each complex number $z_{0}$ a color $(R,G,B)$ where $R,G,B\in[0,1]$ (so that $(1,0,0)$ is red, $(0,1,0)$ is green, $(0,0,1)$ is blue and $(0.5,0,0.5)$ is a light purple, for example). The values of $R$, $G$ and $B$ are determined by looking at the image of said number at each stage of the iteration, $N_{p}^{n}(z_{0})$, and computing the current distance to each of the roots of $p(z)$; so $R=1/d_{r}$ where $d_{r}$ is the positive distance to the root which is colored red, and so on. In this way, the roots themselves are colored red, green, and blue, and every other point gets a mix of the three colors. As the number of iterations increases, each point gets a sharper color, as the sequence of images $\{N_{p}^{n}(z_{0})\}_{n=1}^{\infty}$ converges to one of the three roots. At each stage, the complex plane looks as if out of focus because the coloring function is continuous. As the reader can see in Figure 1, the points at the boundary of each color class form the famous Newton’s fractal (of which, interestingly, Newton was unaware).
@@
Figure 1. Newton’s method approximating $p(z)=z^{3}-2z+2$. Notice the regions of divergence.
Another example of a Newton’s fractal is for $p(z)=z^{3}-1$. The roots of $p(z)$ are the 3rd roots of unity and the Newton map is given by:
$N_{p}(z)=z-\frac{z^{3}-1}{3z^{2}}=\frac{2z^{3}+1}{3z^{2}}$
In this case, there are three basins of attraction (one for each root) and the complement of their union is the Julia set, i.e., the common boundary.
@@
Let $p(z)$ be a non-constant polynomial. $(L_{p},\mathcal{P},\Gamma_{p})$ satisfies the Extendibility Axiom.
@@
[…]
$P_{1}(x)=\frac{2x}{x^{2}+1},$
$P_{2}(x)=\frac{x^{2}-1}{x^{2}+1}.$
Let $\Gamma=\{N_{p}^{n}:n\in\mathbb{N}\}$ where
$N_{p}(x)=\frac{x^{2}+a}{2x}.$
As before, $(L,\mathcal{P},\Gamma)$ is a CCS. Note that $\mathcal{L}=S^{1}$ and that each iterate $N_{p}^{n}$ can be continuously extended to the extended real line $\mathbb{R}\cup\{\infty\}$, i.e., $\mathcal{L}$. For example,
$\tilde{N_{p}}(x)=\begin{cases}\frac{x^{2}+a}{2x},&\text{if }x\in L;\\ \infty,&\text{if }x=0,\infty.\end{cases}$
For every initial guess $x\in L$, the limit $f(x)=\lim_{n\rightarrow\infty}N_{p}^{n}(x)$ converges pointwise to one of the roots. Moreover,
$f(x)=\begin{cases}\sqrt{a},&\text{if }x>0;\\ -\sqrt{a},&\text{if }x<0.\end{cases}$
Notice that $f$ can be extended to $\mathcal{L}$ by
$\tilde{f}(x)=\begin{cases}\sqrt{a},&\text{if }x>0;\\ -\sqrt{a},&\text{if }x<0;\\ \infty,&\text{if }x=0,\infty.\end{cases}$
However, $\tilde{f}$$\mathcal{L}\rightarrow\mathcal{L}$ is not continuous. The set $\overline{\tilde{\Gamma}}$ of deep computations is $\tilde{\Gamma}\cup\{\tilde{f}\}\subseteq B_{1}(\mathcal{L},\mathcal{L})$.
@@
Newton’s method for $p(z)=z^{3}-2z+2$. Let $r_{1}$, $r_{2}$ and $r_{3}$ be the three roots of $p(z)$. Let $B_{1}$, $B_{2}$ and $B_{3}$ be their respective basins. Let $B$ be the basin of the attractive cycle $0,1,0,1,\dots$. Then, $L_{p}=B\cup\bigcup_{i=1}^{3}B_{i}$. Notice that $N_{p}^{n}$ does not converge pointwise. However, the subsequences $N_{p}^{2n}$ and $N_{p}^{2n+1}$ are pointwise convergent to functions $f_{1}$ and $f_{2}$ respectively. $f_{1}$ and $f_{2}$ are two distinct deep computations. Note that for $z\notin L_{p}$, no subsequence of $\tilde{N_{p}^{n}}(z)$ converges to a complex number. However, since $\mathcal{L}=S^{2}$ is compact there is a subsequence of $\tilde{N_{p}^{n}}(z)$ that converges to $\infty$. We can extend $f_{i}:L_{p}\rightarrow\mathcal{L}$ to $\tilde{f_{i}}$$\mathcal{L}\rightarrow\mathcal{L}$ by:
[…]
Again, note that $\tilde{f_{i}}$ for $i=1,2$ are not continuous and that $\tilde{f_{i}}\in\overline{\tilde{\Gamma}}$.
@@ 2.5. Finite precision prefix test
In other words, $\psi_{w}$ determines whether the first $|w|$ bits of a binary sequence is exactly $w$. Let $\Delta=\{\psi_{w}:w\in 2^{<\mathbb{N}}\}$ and $\Gamma$ be the semigroup generated by $\Delta$. Since the sets $\{x\in 2^{\mathbb{N}}:x|_{|w|}=w\}$ are open and closed in $2^{\mathbb{N}}$, then the transitions $\psi_{w}$ are all continuous. In particular, $(L,\mathcal{P},\Gamma)$ satisfies the Extendibility Axiom.
@@ 3.1. NIP, Rosenthal compacta, and deep computations
Let $(L,\mathcal{P},\Gamma)$ be a compositional computational structure (Definition 2.3) satisfying the Extendibility Axiom (Definition 2.5) with $\mathcal{P}$ countable. Let $R$ be an exhaustive collection of sizers. Let $\Delta\subseteq\Gamma$ be $R$-confined. The following are equivalent.
@@
    $\pi_{P}\circ\Delta|_{L[r_{\bullet}]}$ satisfies the NIP for all $P\in\mathcal{P}$ and $r_{\bullet}\in R$; that is, for all $P\in\mathcal{P}$, $r_{\bullet}\in R$, $a<b$, $\{\gamma_{n}\}_{n\in\mathbb{N}}\subseteq\Delta$ there are finite disjoint $E,F\subseteq\mathbb{N}$ such that L[r_]∩⋂_nE(π_Pγ_n)^-1(-,a]∩⋂_nF(π_Pγ_n)^-1[b,)=.
@@
Since $\mathcal{P}$ is countable, $\mathcal{L}[r_{\bullet}]\subseteq\mathbb{R}^{\mathcal{P}}$ is Polish. Also, the Extendibility Axiom implies that $\pi_{P}\circ\tilde{\Delta}|_{\mathcal{L}[r_{\bullet}]}$ is a pointwise bounded set of continuous functions for all $P\in\mathcal{P}$. Hence, Theorem 1.11 and Lemma 1.12 prove the equivalence of (i) and (ii). If (i) holds and $f\in\overline{\Delta}$, then write $f=\mathcal{U}{\rm lim}_{i}\gamma_{i}$ as an ultralimit. Define $\tilde{f}:=\mathcal{U}{\rm lim}_{i}\tilde{\gamma}_{i}$. Hence, for all $r_{\bullet}\in R$ we have $\tilde{f}|_{\mathcal{L}[r_{\bullet}]}\in\overline{\tilde{\Delta}|_{\mathcal{L}[r_{\bullet}]}}\subseteq B_{1}(\mathcal{L}[r_{\bullet}],\mathcal{L}[r_{\bullet}])$. That every deep computation is a pointwise limit of a countable sequence of computations follows from the fact that for a Polish space $X$ every compact subset of $B_{1}(X)$ is Fréchet-Urysohn (that is, a space where topological closures coincide with sequential closures, see Theorem 3F in [BFT78] or Theorem 4.1 in [DEB13]). ∎

### v4

@@ 2.4. Newton’s method as a CCS
We formalize the Newton-Raphson root-finding method as a CCS [GAL00], [BFB16, Section 2.3].
2.4.1. The extended complex plane
Let $\hat{\mathbb{C}}$$\mathbb{C}\cup\{\infty\}$ be the extended complex plane (the Alexandroff —one-point— compactification of $\mathbb{C}$). We regard the unit sphere $S^{2}\subseteq\mathbb{R}^{3}$ as the Riemann sphere and endowed with the stereographic projection map $S^{2}\to\hat{\mathbb{C}}$ —a topological homeomorphism whose inverse $\hat{\mathbb{C}}\rightarrow S^{2}$ is given by $z\mapsto(P_{1}(z),P_{2}(z),P_{3}(z))$, where $(P_{1}(\infty),P_{2}(\infty),P_{3}(\infty))=(0,0,1)$ and, for $z\neq\infty$,
$P_{1}(z)$ $=\frac{2{\Re}(z)}{|z|^{2}+1},$ $P_{2}(z)$ $=\frac{2{\Im}(z)}{|z|^{2}+1},$ $P_{3}(z)$ $=\frac{|z|^{2}-1}{|z|^{2}+1}.$
The extended complex plane $\hat{\mathbb{C}}$ is thus endowed with the real predicates $P_{1},P_{2},P_{3}$ with values in [1,1][-1,1]; its topology is metrized by the pullback of the usual Euclidean metric on $S^{2}$.
2.4.2. Newton maps
Let $p=p(z)$ be a non-constant polynomial with complex coefficients. [^8] The entire discussion of Newton CCSs applies, mutatis mutandis, to rational maps $p$. The Newton method step (or Newton map) for $p$ is the map
(2.7) $N_{p}:z\mapsto z-\frac{p(z)}{p^{\prime}(z)}.$
Now, if one wishes to interpret such map as a (complex-valued) function of a complex variable, its natural domain $\mathbb{C}\setminus\mathcal{Z}(p^{\prime})$ excludes all roots of $p^{\prime}$. If one wishes to iterate the Newton map indefinitely, the natural domain is
$L_{p}\coloneqq\mathbb{C}\setminus\bigcup_{n\in\mathbb{N}}N_{p}^{-n}[\mathcal{Z}(p^{\prime})]=\{z\in\mathbb{C}:\text{$N_{p}^{(n)}(z)\notin\mathcal{Z}(p^{\prime})$ for all $n\in\mathbb{N}$}\}.$
On the other hand, any rational map extends continuously to an analytic (meromorphic) function on the extended complex plane, so $N_{p}$ naturally defines a map $\hat{\mathbb{C}}\to\hat{\mathbb{C}}$.
Successive iterations of the Newton-Raphson method (i.e., compositional powers $N_{p}^{(n)}$) may be regarded as computations of a suitable CCS whose states space is perhaps $L_{p}\subseteq\mathbb{C}$, or perhaps the full extended complex plane $\hat{\mathbb{C}}$. In the former case, it is natural to use the predicates $X:z\mapsto\Re(z)$ and $Y:z\mapsto\Im z$; in the latter, the predicates $P_{1},P_{2},P_{3}$ are necessary because $X,Y$ are not defined at $\infty$.
An important (if somewhat obvious) point is that the choice of predicates on a CCS greatly affect the topologies of itself and of its type space. Thus, the type space of $(L_{p},\{X,Y\})$ above is $\mathbb{R}^{2}$ ($\cong\mathbb{C}$), while the type space of $(\hat{\mathbb{C}},\{P_{1},P_{2},P_{3}\})$ is $S^{2}$ $\cong\hat{\mathbb{C}}$. All $\{P_{1},P_{2},P_{3}\}$-types are realized in $\hat{\mathbb{C}}$, but not all $\{X,Y\}$-types are realized in $L_{p}$ (let alone all $\{P_{1},P_{2},P_{3}\}$-types).
2.4.3. Newton CCSs
Per the discusion above, we choose to discuss the Newton-Raphson method as a CCS whose states space is the extended complex plane $\hat{\mathbb{C}}$.
@@ 2.4.3. Newton CCSs
Let $p=p(z)$ be a nonconstant complex polynomial. We say that $(\hat{\mathbb{C}},\mathcal{P},\Gamma_{p})$ is Newton’s method for $p(z)$ if:
  • $\mathcal{P}:=\{P_{1},P_{2},P_{3}\}$ are the coordinates of inverse steregraphic projection (as above);
  • $\Gamma_{p}:=\{N_{p}^{(n)}:n\in\mathbb{N}\}$ consists of iterates of the Newton map $N_{p}$ in equation (2.7).
Given an initial estimate $z_{0}$, one hopes for the sequence $N_{p}^{(\bullet)}(z_{0})\coloneqq\bigl(N_{p}^{(n)}(z_{0})\bigr)_{n}$ of Newton steps applied to $z_{0}$ to converge to a root $r$ of $p$. However, Newton’s method fails to converge to any root in many cases. For example, consider the polynomial $p(z)=z^{3}-2z+2$ with Newton map
$N_{p}(z)=z-\frac{z^{3}-2z+2}{3z^{2}-2}=\frac{2z^{3}-2}{3z^{2}-2}.$
[…]
For any $n\in\mathbb{N}$ we define a coloring of $\mathbb{C}$ in the natural way: color each $z_{0}\in\mathbb{C}$ using the standard color of $z_{n}\coloneqq N_{p}^{(n)}(z_{0})$. At any stage, particularly early ones, the coloring looks “fuzzy” (out of focus) because the coloring function is continuous; however, as $n$ increases, the coloring becomes sharper. For one thing, for many $z$ in the Fatou set, the sequence $(z_{n})$ converges to one of the three roots; the colors of $(z_{n})$ tends towards (a single one of) red, green or blue correspondingly. As immediately recognizable in Figure 1, such $z_{0}$ lie in large (open connected) regions. However, the Fatou set also includes points $z_{0}$ for which $(z_{n})$ diverges, but has some subsequence converging uniformly in some compact neighborhood of $z_{0}$; those points form dark teal regions. The Julia set at boundary of the Fatou set is Newton’s fractal proper.
@@
Figure 1. Behavior of iterates of the Newton map for $p(z)=z^{3}-2z+2$. In (C), the “color fuzzines” of (A) and (B) has already visually disappeared; only sharply defined open red, green and blue regions (of convergence towards a specific root), as well as dark teal open regions (of divergence) remain visually identifiable: their union (the Fatou set of $p$) is dense in the limit, and the Newton fractal of $p$ is the closed nowhere dense complement (the Julia set of $p$).
Consider next the $3$-cyclotomic polynomial $p(z)=z^{3}-1$. The roots of $p(z)$ are the cube roots of unity; the Newton map is
$N_{p}:z\mapsto z-\frac{z^{3}-1}{3z^{2}}=\frac{2z^{3}+1}{3z^{2}}.$
In Figure 2(C), one sees a Fatou set consisting of solid-color (disconnected) open sets, one in each of the colors red (for the root $r_{0}=1$), green and blue (for $r_{\pm 1}=\exp(\pm 2\pi i/3)$, resp.), where the iterates sequence $(z_{n})=(p^{(n)}(z_{0}))_{n}$ starting with any colored point $z_{0}$ converges to the corresponding root. (Thus, each solid-color region is the basin of attraction of the respective root).
@@
If $p(z)$ is a non-constant polynomial, then $(L_{p},\mathcal{P},\Gamma_{p})$ satisfies the Extendibility Axiom.
@@
Computation of square roots. Let $a$ be a positive real number and $p(x)=x^{2}-a$. Let $\mathbb{R}\to S^{1}:x\mapsto(P_{1}(x),P_{2}(x))$ be the (inverse) stereographic projection of the real line onto $S^{1}\subseteq\mathbb{R}^{2}$, i.e.,
$P_{1}(x)$ $=\frac{2x}{x^{2}+1},$ $P_{2}(x)$ $=\frac{x^{2}-1}{x^{2}+1},$
extended to a projective transformation of $\mathbb{R}\mathbb{P}^{1}=\mathbb{R}\cup\{\infty\}$ by $\infty\mapsto(P_{1}(\infty),P_{2}(\infty))=(0,1)$. Let $\mathcal{P}=(P_{1},P_{2})$ and $\Gamma=\{N_{p}^{n}:n\in\mathbb{N}\}$, where $N_{p}:\mathbb{R}\mathbb{P}^{1}\rightarrow\mathbb{R}\mathbb{P}^{1}$ is given by
$x\mapsto x-\frac{x^{2}-a}{2x}=\begin{cases}\frac{x^{2}+a}{2x}&(x\neq 0,\infty),\\ \infty&(x=0,\infty).\end{cases}$
As before, $(L,\mathcal{P},\Gamma)$ is a CCS with $L\coloneqq\mathbb{R}\mathbb{P}^{1}\setminus\{0,\infty\}=\mathbb{R}\setminus\{0\}$ whose type space is $\mathcal{L}=S^{1}\cong\mathbb{R}\mathbb{P}^{1}$. In fact (with the identification $\mathcal{L}\cong\mathbb{R}\mathbb{P}^{1}$), the (full) sequence if iterates of the Newton map has a pointwise limit $\mathfrak{f}:\mathcal{L}\to\mathcal{L}$, namely
$\mathfrak{f}(x)=\lim_{n\rightarrow\infty}N_{p}^{n}(x)=\begin{cases}\sqrt{a}&(x>0),\\ -\sqrt{a}&(x<0),\\ \infty&(x=0,\infty).\end{cases}$
Thus, $\mathfrak{f}$ is discontinuous at $x=0,\infty$. The set of deep computations is $\{N_{p}^{n}\}_{n}\cup\{\mathfrak{f}\}\subseteq B_{1}(\mathcal{L},\mathcal{L})$.
@@
Newton’s method for $p(z)=z^{3}-2z+2$. Let $B_{i}$, $B_{\pm 1}$ be the basin of the root $r_{i}$ ($i=0,\pm 1$), and let $B$ be the basin of the attractive cycle 0, 1, 0, 1, …. It may be shown that the Fatou set of $N_{p}$ is $F_{p}=B\cup\bigcup_{i=0,\pm 1}B_{i}$, each of these four sets being nonempty open. Since $B$ is nonempty, $N_{p}^{n}$ does not converge pointwise; however, the subsequences $N_{p}^{2n}$ and $N_{p}^{2n+1}$ are pointwise convergent to two distinct deep computations $f_{0}$ (constant 0 on $B$) and $f_{1}$ (constant $1$ on BB). [^10] As remarked above, $\infty$ is a repelling fixed point of $N_{p}$, hence $f_{0}(\infty)=\infty=f_{1}(\infty)$. The behavior of the iterates $N_{p}$ on the Julia set $\mathbb{C}\setminus L_{p}$ is chaotic
@@ 2.6. Finite precision prefix test
In other words, $\psi_{w}$ determines whether the first $|w|$ bits of a binary sequence is exactly $w$. Let $\Delta=\{\psi_{w}:w\in 2^{<\mathbb{N}}\}$ and $\Gamma$ be the semigroup generated by $\Delta$. Since the sets $\{x\in 2^{\mathbb{N}}:x|_{|w|}=w\}$ are open and closed in $2^{\mathbb{N}}$, we see that the transitions $\psi_{w}$ are all continuous. In particular, $(L,\mathcal{P},\Gamma)$ satisfies the Extendibility Axiom.
@@ 3.1. NIP, Rosenthal compacta, and deep computations
Let $(L,\mathcal{P},\Gamma)$ be a compositional computational structure (Definition 2.4) satisfying the Extendibility Axiom (Definition LABEL:D:extendibility_axiom) with $\mathcal{P}$ countable. Let $R$ be an exhaustive collection of sizers. Let $\Delta\subseteq\Gamma$ be $R$-confined. The following are equivalent.
@@
    $\pi_{P}\circ\Delta|_{L[r_{\bullet}]}$ satisfies the NIP for all $P\in\mathcal{P}$ and $r_{\bullet}\in R$; that is, for all $P\in\mathcal{P}$, $r_{\bullet}\in R$, $a<b$, $\{\gamma_{n}\}_{n\in\mathbb{N}}\subseteq\Delta$ there are finite disjoint $E,F\subseteq\mathbb{N}$ such that
    $L[r_{\bullet}]\cap\bigcap_{n\in E}(\pi_{P}\circ\gamma_{n})^{-1}(-\infty,a]\cap\bigcap_{n\in F}(\pi_{P}\circ\gamma_{n})^{-1}[b,\infty)=\emptyset.$
@@
Since $\mathcal{P}$ is countable, $\mathcal{L}[r_{\bullet}]\subseteq\mathbb{R}^{\mathcal{P}}$ is Polish. Also, the Extendibility Axiom implies that $\pi_{P}\circ\tilde{\Delta}|_{\mathcal{L}[r_{\bullet}]}$ is a pointwise bounded set of continuous functions for all $P\in\mathcal{P}$. Hence, Theorem 1.11 and Lemma 1.12 prove the equivalence of (i) and (ii). If (i) holds and $f\in\overline{\Delta}$, then write $f=\mathcal{U}{\rm lim}_{i}\gamma_{i}$ as an ultralimit. Define $\tilde{f}:=\mathcal{U}{\rm lim}_{i}\tilde{\gamma}_{i}$. Hence, for all $r_{\bullet}\in R$ we have $\tilde{f}|_{\mathcal{L}[r_{\bullet}]}\in\overline{\tilde{\Delta}|_{\mathcal{L}[r_{\bullet}]}}\subseteq B_{1}(\mathcal{L}[r_{\bullet}],\mathcal{L}[r_{\bullet}])$. That every deep computation is a pointwise limit of a countable sequence of computations follows from the fact that for a Polish space $X$ every compact subset of $B_{1}(X)$ is Fréchet-Urysohn (that is, a space where any point in the closure of a set $A$ is the limit of a sequence of points in $A$; see Theorem 3F in [BFT78] or Theorem 4.1 in [DEB13]). ∎